# web_appendix

### v4

@@ Abstract
We propose difference-in-differences estimators in designs where the treatment is continuously distributed at every period, as is often the case when one studies the effects of taxes, tariffs, or prices. We assume that between consecutive periods, the treatment of some units, the switchers, changes, while the treatment of other units remains constant. We show that under a placebo-testable parallel-trends assumption, averages of the slopes of switchers’ potential outcomes can be nonparametrically estimated. We generalize our estimators to the instrumental-variable case. We use our estimators to estimate the price-elasticity of gasoline consumption.
Keywords: differences-in-differences, continuous treatment, two-way fixed effects regressions, heterogeneous treatment effects, panel data, policy evaluation, instrumental variable.
JEL Codes: C21, C23
@@ 1 Introduction
A popular method to estimate the effect of a treatment on an outcome is to estimate a two-way fixed effects (TWFE) regression that controls for unit and time fixed effects:
@@
where $D_{i,t}$ is the treatment of unit $i$ at time $t$. de Chaisemartin and D’Haultfœuille (2023a) find that 26 of the 100 most cited papers published by the American Economic Review from 2015 to 2019 have estimated at least one TWFE regression. de Chaisemartin and D’Haultfœuille (2020), Goodman-Bacon (2021), and Borusyak et al. (2021) have shown that under a parallel trends assumption, TWFE regressions are not robust to heterogeneous effects: they may estimate a weighted sum of treatment effects across periods and units, with some negative weights. Owing to the negative weights, $\beta_{TWFE}$ could be, say, negative, even if the treatment effect is positive for every unit $\times$ period. Importantly, the result in de Chaisemartin and D’Haultfœuille (2020) applies to binary, discrete, and continuous treatments.
Several alternative heterogeneity-robust difference-in-difference (DID) estimators have been proposed (see Table 2 of de Chaisemartin and D’Haultfœuille, 2023b). Some apply to binary and staggered treatments (see Sun and Abraham, 2021; Callaway and Sant’Anna, 2021; Borusyak et al., 2021). Some apply to designs where all units start with a treatment equal to 00, and then get treated with heterogeneous, potentially continuously distributed treatment intensities (see de Chaisemartin and D’Haultfœuille, 2023a; Callaway et al., 2021; de Chaisemartin and D’Haultfœuille, 2024). However, treatments continuously distributed at every period, including the first one in the data, are ubiquitous in applied work. For instance, taxes (see Li et al., 2014) or tariffs (see Fajgelbaum et al., 2020) are often continuously distributed at all periods. No difference-in-difference estimator robust to heterogeneous effects is available for such designs. Proposing such estimators is the purpose of this paper.
We assume that we have a panel data set, whose units could be geographical locations such as counties. We start by considering the case where the panel has two time periods. From period one to two, the treatment of some units, hereafter referred to as the switchers, changes. On the other hand, the treatment of other units, hereafter referred to as the stayers, does not change. We consider two target parameters. The first one is the average slope of switchers’ period-two potential outcome function, from their period-one to their period-two treatment, hereafter referred to as the Average of Slopes (AS). Our second target is a weighted average of switchers’ slopes, where switchers receive a weight proportional to the absolute value of their treatment change, hereafter referred to as the Weighted Average of Slopes (WAS). We propose a novel parallel trends assumption on the outcome evolution of switchers and stayers with the same period-one treatment, in the counterfactual where switchers’ treatment would not have changed. Because it conditions on units’ period-one treatment, this parallel-trends assumption does not impose any restriction on effect heterogeneity. This parallel trends assumption is also placebo testable, by comparing switchers’ and stayers’ outcome evolutions before switchers’ treatment changes. We view the possibility of placebo-testing it as an important advantage of our assumption, as placebo tests are an essential step in establishing the credibility of an identifying assumption in observational studies (Imbens et al., 2001; Imbens and Xu, 2024). We show that under our placebo-testable parallel-trends assumption, the AS and the WAS are identified. This contrasts with other target parameters, like the dose-response function or the average marginal effect, which can only be identified under non-placebo-testable assumptions.
Economically, the AS and the WAS can serve different purposes, so neither parameter dominates the other. Under shape restrictions on the potential outcome function, the AS can be used to infer the effect of other treatment changes than those that took place from period one to two. Instead, the WAS can be used to conduct a cost-benefit analysis of the treatment changes that effectively took place. On the other hand, when it comes to estimation, the WAS unambiguously dominates the AS. First, we show that it can be estimated at the standard parametric rate even if switchers can experience an arbitrarily small change of their treatment between consecutive periods. Second, we show that under some conditions, the asymptotic variance of the WAS estimator is strictly lower than that of the AS estimator. Third, unlike the AS, the WAS is amenable to doubly-robust estimation, which comes with a number of advantages.
@@
Then, we consider the instrumental-variable (IV) case. For instance, one may be interested in estimating the price-elasticity of a good’s consumption. If prices respond to demand shocks, the counterfactual consumption trends of units experiencing and not experiencing a price change may not be parallel. On the other hand, the counterfactual consumption trends of units experiencing and not experiencing a tax change may be parallel. Then, taxes may be used as an instrument for prices. In such cases, we show that the reduced-form WAS effect of the instrument on the outcome divided by the first-stage WAS effect is equal to a weighted average of switchers’ outcome-slope with respect to the treatment, where switchers with a larger first-stage effect receive more weight. Hereafter, we refer to this effect as the IV-WAS effect. The ratio of the reduced-form and first-stage AS effects is also equal to a weighted average of slopes, with arguably less natural weights, so in the IV case the WAS seems both economically and statistically preferable to the AS. Importantly, we show that the reduced-form parallel-trends assumption implicitly restricts treatment-effect heterogeneity. Such restrictions can be alleviated by controlling for groups baseline treatment in the IV specification, but the resulting estimator still restricts effects’ heterogeneity across units.
We consider other extensions. First, we extend our results to applications with more than two time periods. Importantly, with several time periods our estimators rely on a parallel-trends assumption over consecutive periods, rather than over the entire duration of the panel. Second, we propose a placebo estimator of the parallel-trends assumption underlying our estimators.
@@
Finally, we use the yearly, 1966 to 2008 US state-level panel dataset of Li et al. (2014) to estimate the effect of gasoline taxes on gasoline consumption and prices. Using the WAS estimators, we find a significantly negative effect of taxes on gasoline consumption, and a significantly positive effect on prices. The AS estimators are close to, and not significantly different from, the WAS estimators, but they are also markedly less precise: their standard errors are almost three times larger than that of the WAS estimators. This shows that our theoretical result on the precision ranking of the WAS and AS estimators, which we derive under strong assumptions, still holds in a real-life application where those assumptions probably do not hold. The precision losses attached to using the AS have consequences. The AS estimator of the effect of taxes on prices is not statistically significant, so with that estimator one cannot use taxes as an instrument to estimate the price-elasticity of consumption, because the instrument does not have a first-stage. This contrasts with the WAS, whose first-stage t-stat is around 7. We compute an IV-WAS estimator of the price elasticity of gasoline consumption, and find a fairly small elasticity of -0.67, in line with previous literature (for instance, Hausman and Newey, 1995, find a long-run elasticity of -0.81). Our placebo estimators are small, insignificant, and fairly precisely estimated, thus suggesting that our parallel trends assumption is plausible.
Our estimators are computed by the did_multiplegt_stat Stata package, available from the SSC repository. Our package allows estimators with control variables and weights, see the help file and the package’s companion paper for further details.
@@
Our paper builds upon several previous papers in the panel data literature. Chamberlain (1982) seems to be the first paper to have proposed an estimator of the AS parameter. Under the assumption of no counterfactual time trend, the estimator therein is a before-after estimator. Then, our paper is closely related to the work of Graham and Powell (2012), who also propose DID estimators of the AS (see their Equation (21)) when the treatment is continuously distributed at every time period. Their estimators rely on a linear effect assumption and assume that units experience the same evolution of their treatment effect over time, a parallel-trends-on-treatment-effects assumption. By contrast, our estimator of the AS does not place any restriction on treatment effects. But our main contribution to this literature is to introduce the WAS, and to contrast the pros and cons of the AS and WAS estimators. Our results are also related to Hoderlein and White (2012), who consider the average marginal effect of a continuous treatment with panel data. However, their target parameters and identifying assumptions are different. For instance, they rule out systematic changes of the outcome over time.
@@
With respect to the aforementioned heterogeneity-robust DID literature, we make two contributions. First, in the non-IV case we propose estimators that can be used even if units’ treatment varies at baseline. Thus we usefully complement previous literature, that has mostly focused on the case where all units have a baseline treatment equal to zero (see Sun and Abraham, 2021; Callaway and Sant’Anna, 2021; Borusyak et al., 2021; Callaway et al., 2021; de Chaisemartin and D’Haultfœuille, 2024). One exception predating this paper is de Chaisemartin and D’Haultfœuille (2020), who allow for a non-binary discrete treatment at baseline in their Web Appendix, and propose estimators comparing switchers and stayers with the same baseline treatment. However, that paper does not allow for continuously distributed treatments, and comparing switchers and stayers with the same baseline treatment is no longer feasible with a continuously distributed treatment. [^1] Another exception, posterior to this paper, is de Chaisemartin and D’Haultfœuille (2023a), who extend the estimators in this paper to models with dynamic effects in Section 1.10 of their Web Appendix. Second, in the IV case, previous IV-DID literature has only considered classical designs with two periods and binary instrument and treatment (de Chaisemartin, 2010; Hudson et al., 2017), as well as fuzzy DID designs, a special case of IV-DIDs (de Chaisemartin and D’Haultfœuille, 2018). Instead, this paper proposes broadly applicable IV-DID estimators, and also highlights that IV-DID estimators impose restrictions on treatment effect-heterogeneity, which can be mitigated by controlling for the baseline treatment.
Importantly, our estimators require that there be some stayers, whose treatment does not change between consecutive periods. This assumption is unlikely to be met when the treatment is say, precipitations: for instance, US counties never experience the exact same precipitations over two consecutive years. In de Chaisemartin et al. (2023), we discuss the (non-trivial) extension of the results in this paper to applications without stayers.
@@
In Section 2, we present the set-up, introduce notation and discuss our main assumptions. In Section 3, we introduce the AS and discuss its identification and estimation. Section 4 then turns to the WAS. Section 5 extends our previous results to an instrumental variable set-up. We consider other extensions in Section 6. Finally, our application is developed in Section 7. The proofs are collected in the appendix.
@@ 2.1 Set-up
A representative unit is drawn from an infinite super population, and observed at two time periods. This unit could be an individual or a firm, but it could also be a geographical unit, like a county or a region. [^2] In that case, one may want to weight the estimation by counties’ or regions’ populations. Extending the estimators we propose to allow for such weighting is a mechanical extension. All expectations below are taken with respect to the distribution of variables in the super population. We are interested in the effect of a continuous and scalar treatment variable on that unit’s outcome. Let $D_{1}$ (resp. $D_{2}$) denote the unit’s treatment at period 1 (resp. 2), and let $\mathcal{D}_{1}$ (resp. $\mathcal{D}_{2}$) be the set of values $D_{1}$ (resp. $D_{2}$) can take, i.e. its support. Let $S=1\{D_{2}\neq D_{1}\}$ be an indicator equal to 1 if the unit’s treatment changes from period one to two, i.e. if they are a switcher.
For any $d\in\mathcal{D}_{1}\cup\mathcal{D}_{2}$, let $Y_{1}(d)$ and $Y_{2}(d)$ respectively denote the unit’s potential outcomes at periods 1 and 2 with treatment $d$, and let $Y_{1}$ and $Y_{2}$ denote their observed outcomes at periods 1 and 2. Our potential outcome notation assumes that $Y_{1}$ does not depend on units’ period-two treatment, thus ruling out anticipation effects, a commonly-made assumption in the DID literature. Our notation also assumes that $Y_{2}$ does not depend on units’ period-one treatment, thus ruling out dynamic effects. When the treatment is continuously distributed at period one, allowing for dynamic effects opens up the so-called initial-conditions problem. As units receive heterogeneous doses at period one, they may have experienced treatment changes before period one. With dynamic effects such changes may still affect their outcome over the study period, but they cannot be accounted for because they are not observed. Ruling out dynamic effects allows us to abstract from this thorny issue, but could yield misleading results if dynamic effects are present. To alleviate this concern, in Section 6.3 we propose a modified version of our estimators, robust to dynamic effects up to a pre-specified treatment lag.
@@ 2.2 Assumptions
We make the following assumptions. [^3] Throughout the paper, we implicitly assume that all potential outcomes have an expectation.
@@
(Parallel trends) For all $d_{1}\in\mathcal{D}_{1}$, $E(\Delta Y(d_{1})|D_{1}=d_{1},D_{2})=E(\Delta Y(d_{1})|D_{1}=d_{1})$.
Assumption 1 is a parallel trends assumption, requiring that $\Delta Y(d_{1})$ be mean independent of $D_{2}$, conditional on $D_{1}=d_{1}$.
@@
(Bounded treatment, Lipschitz and bounded potential outcomes)
@@
    For all $t\in\{1,2\}$ and for all $(d,d^{\prime})\in\mathcal{D}_{t}^{2}$, there is a random variable $\overline{Y}\geq 0$ such that $|Y_{t}(d)-Y_{t}(d^{\prime})|\leq\overline{Y}|d-d^{\prime}|$, with $\sup_{(d_{1},d_{2})\in\text{Supp}(D_{1},D_{2})}E[\overline{Y}|D_{1}=d_{1},D_{2}=d_{2}]<\infty$.
Assumption 2 ensures that all the expectations below are well defined. It requires that the set of values that the period-one and period-two treatments can take be bounded. It also requires that the potential outcome functions be Lipschitz (with a unit-specific Lipschitz constant). This will automatically hold if $d\mapsto Y_{2}(d)$ is differentiable with respect to $d$ and has a bounded derivative.
@@
(iid sample) We observe $(Y_{i,1},Y_{i,2},D_{i,1},D_{i,2})_{1\leq i\leq n}$, that are independent and identically distributed vectors with the same probability distribution as $(Y_{1},Y_{2},D_{1},D_{2})$.

### v5

@@ Abstract
We propose difference-in-differences (DID) estimators in designs where the treatment is continuously distributed in every period, as is often the case when one studies the effects of taxes, tariffs, or prices. We assume that between consecutive periods, the treatment of some units, the switchers, changes, while the treatment of other units, the stayers, remains constant. We show that under a parallel-trends assumption, weighted averages of the slopes of switchers’ potential outcomes are nonparametrically identified by difference-in-differences estimands comparing the outcome evolutions of switchers and stayers with the same baseline treatment. Controlling for the baseline treatment ensures that our estimands remain valid if the treatment’s effect changes over time. We highlight two possible ways of weighting switcher’s slopes, and discuss their respective advantages. For each weighted average of slopes, we propose a doubly-robust, nonparametric, $\sqrt{n}$-consistent, and asymptotically normal estimator. We generalize our results to the instrumental-variable case. Finally, we apply our method to estimate the price-elasticity of gasoline consumption.
Keywords: differences-in-differences, continuous treatment, instrumental variable.
@@ 1 Introduction
To estimate a treatment’s effect, researchers often run a two-way fixed effects (TWFE) regression:
@@
where $D_{i,t}$ is the treatment of unit $i$ at time $t$, and $\alpha_{i}$ and $\gamma_{t}$ are unit and period fixed effects. de Chaisemartin and D’Haultfœuille (2024) find that 26 of the 100 most cited papers published by the American Economic Review from 2015 to 2019 estimate TWFE regressions. Chiu et al. (2024) find that 52 papers published by the top three political science journals from 2017 to 2022 estimate a TWFE regression. Researchers have long believed that those regressions only rely on a parallel-trends assumption, requiring that without treatment, all units would have experienced the same outcome evolutions. However, recent papers have shown that under that assumption alone, TWFE regressions may estimate a non-convex sum of treatment effects across units and periods (see, e.g., de Chaisemartin and D’Haultfœuille, 2020; Borusyak et al., 2024). Then, $\beta_{TWFE}$ could be, say, negative, even if the treatment effect is positive for every $(i,t)$.
In reaction to this finding, several “heterogeneity-robust difference-in-difference” estimators, consistent for convex combinations of $(i,t)$-specific effects under a parallel-trends assumption, have been proposed. However, none can be applied to treatments continuously distributed at every period. Yet, such treatments are commonly found. In economics, taxes (see Li et al., 2014) or tariffs (see Fajgelbaum et al., 2020) are often continuously distributed at all periods. Similarly, Peterson (2021), Aklin and Kern (2021), and McNamara et al. (2022) are examples of papers in political sciences and epidemiology studying treatments continuous in all periods. With continuous treatments, due to the lack of an untreated group, TWFE regressions often estimate a highly non-convex combination of effects. For instance, with two time periods, exactly 50% of $(i,t)$-specific effects are weighted negatively (de Chaisemartin and Lei, 2021). Then, proposing robust estimators for such treatments is important.
We assume that we have a two-periods panel data set. From period one to two, the treatment of some units, referred to as the switchers, changes, while the treatment of some other units, the stayers, does not change. We propose a novel parallel-trends assumption, which requires that switchers and stayers with the same period-one treatment would have experienced the same average outcome evolutions from period one to two, if switchers’ treatment had not changed. Our assumption has two desirable features. First, because it conditions on units’ period-one treatment, it does not restrict effects’ heterogeneity. Instead, as explained later, assuming parallel trends for switchers and stayers with different period-one treatments would essentially amount to assuming that the treatment effect is constant over time. Second, when another prior period of data, period zero, is available, the plausibility of our assumption can be assessed via a “pre-trends” or “placebo” test, comparing the period-zero-to-one outcome evolutions of period-one-to-two switchers and stayers. [^1] As explained later, the pre-trends estimator has to be computed in the subsample of units whose treatment does not change from period zero to one. Then, one may also restrict the main estimators to that subsample. Thus, one can assess if switchers and stayers are on parallel trends before switchers’ treatment changes.
Then, we consider two target parameters. The first is the average slope of switchers’ period-two potential outcome function, from their period-one to their period-two treatment, referred to as the Average of Slopes (AS). The second is a weighted average of switchers’ slopes, where switchers receive a weight proportional to the absolute value of their treatment change, referred to as the Weighted Average of Slopes (WAS). The AS and WAS serve different purposes. Under shape restrictions on the potential outcome function, the AS can be used to identify or bound the effect of other treatment changes than those that took place. Instead, the WAS can be used to conduct a cost-benefit analysis of the treatment changes that took place. Under our parallel-trends assumption, whose plausibility can be assessed using a pre-trends test, the AS and the WAS are identified by DID estimands that compare the outcome evolution of switchers and stayers with the same period-one treatment. This is in contrast with other targets considered in the literature like the dose-response function or the average marginal effect, which can only be identified under assumptions whose plausibility cannot be assessed via a pre-trends test.
Turning to estimation, the challenge with a continuous treatment is that the sample does not contain switchers and stayers with the same period-one treatment. In other words, the estimands identifying the AS and WAS depend on nonparametric regression functions. Therefore, we derive doubly-robust moment conditions identifying the AS and WAS, and we propose doubly-robust estimators. We show that the AS estimator converges at the parametric rate, provided switchers cannot experience arbitrarily small treatment changes. We then show that the WAS can always be estimated at the parametric rate. Under some conditions, the asymptotic variance of the WAS estimator is strictly lower than that of the AS estimator.
@@
We extend those results in several directions. First, we consider the instrumental-variable (IV) case, where one makes a parallel-trends assumption with respect to an instrument rather than the treatment. For instance, one may be interested in estimating the price-elasticity of a good. If prices respond to demand shocks, the counterfactual consumption trends of units experiencing and not experiencing a price change may not be the same, so a parallel-trends assumption with respect to prices may not hold. But instead, one can make a parallel-trends assumption with respect to taxes, used as an instrument for prices. We start by noting that parallel-trends with respect to an instrument restricts treatment-effect heterogeneity. Then, we show that under that assumption, the reduced-form WAS effect of the instrument on the outcome divided by the first-stage WAS effect is equal to a weighted average of switchers’ outcome-slope with respect to the treatment, where switchers with a larger first-stage effect receive more weight.
We also consider other extensions. First, we consider applications with more than two time periods. Second, we propose a pre-trends estimator to assess the plausibility of our parallel-trends assumption. Third, with more than two time periods, our estimators rule out dynamic effects: they assume that units’ outcome at period $t$ only depends on their period-$t$ treatment, not on their prior treatments. To alleviate this concern, we propose a modified version of our estimators, robust to dynamic effects up to a pre-specified treatment lag. Finally, we propose estimators allowing for control variables. The did_multiplegt_stat Stata package, under development but already available from the SSC repository, computes some of our estimators.
@@
As an illustration, we use the yearly, 1966 to 2008 US state-level panel dataset of Li et al. (2014) to estimate the effect of gasoline taxes on gasoline consumption and prices. Using the WAS estimators, we find a significantly negative effect of taxes on gasoline consumption, and a significantly positive effect on prices. The AS estimators are close to, and not significantly different from, the WAS estimators, but they are insignificant because they are markedly less precise: their standard errors are almost three times larger. We compute an IV estimator of the price elasticity of gasoline consumption, and find a significantly negative elasticity of -0.66, which is more negative than, but not significantly different from, that reported in recent literature (see, e.g., Coglianese et al., 2017). Our pre-trend estimators are small and insignificant, thus lending credibility to our parallel-trends assumption.
@@
Our paper builds upon several papers in the panel data literature. Chamberlain (1982) seems to be the first paper to have proposed an estimator of the AS parameter. Under the assumption of no counterfactual time trend, the estimator therein is a before-after estimator. Our paper is also closely related to the work of Graham and Powell (2012), who propose DID estimators of the AS (see their Equation (21)) when the treatment is continuously distributed at every time period. Their estimators rely on a linear effect assumption and assume that units experience the same evolution of their treatment effect over time. By contrast, our estimator of the AS does not place any restriction on treatment effects. But our main contribution to this literature is to introduce the WAS, and to contrast the different purposes served by the two parameters.
@@
With respect to the aforementioned heterogeneity-robust DID literature, we make two contributions. First, in the non-IV case we propose estimators that can be used even if units receive heterogeneous treatment doses at baseline. Thus we complement previous literature, that has mostly focused on the case where all units receive the same dose at baseline (see Sun and Abraham, 2021; Callaway and Sant’Anna, 2021; Liu et al., 2024; Borusyak et al., 2024; de Chaisemartin et al., 2024; Callaway et al., 2021). One exception predating this paper is de Chaisemartin and D’Haultfœuille (2020), who, in their web appendix, consider designs where units receive heterogeneous doses of a non-binary discrete treatment at baseline. With a discrete treatment, they can propose DID estimators comparing switchers and stayers with the exact same baseline treatment. This solution is no longer applicable with a continuously distributed treatment, hence the estimation challenge we tackle in this paper. Another exception, contemporaneous of this paper, is de Chaisemartin and D’Haultfœuille (2024). They propose estimators robust to dynamic effects up to any lag, focusing mostly on discrete treatments. In Section 1.10 of their web appendix, they build upon the current paper to sketch an extension of their estimators to continuous treatments. However, contrary to here, the estimators therein are parametric and not doubly-robust, and the estimators’ asymptotic distribution is not derived. Moreover, they show that with dynamic effects, under a parallel-trends assumption one can identify weighted averages of potential outcome’s slopes with respect to various treatment lags. Those effects are harder to interpret than the effects identified in this paper.
In the IV case, we extend de Chaisemartin (2010) and Hudson et al. (2017), who consider designs with a binary instrument and treatment. Our paper is also the first to note that parallel-trends with respect to an instrument restricts treatment-effect heterogeneity.
@@
Section 2 presents the set-up, our main assumptions, and a building-block identification result. Section 3 considers identification and estimation of the AS. Section 4 turns to the WAS. Section 5 sketches some extensions, covered in our web appendix. Finally, Section 6 presents our application. The web appendix also collects all proofs.
@@ 2.1 Set-up
A representative unit is drawn from an infinite super population, and observed at two time periods. This unit could be an individual or a firm, but it could also be a geographical unit, like a county or a region. [^2] In that case, one may want to weight the estimation by counties’ or regions’ populations. Extending the estimators we propose to allow for such weighting is a mechanical extension. All expectations below are taken with respect to the distribution of variables in the super population. We are interested in the effect of a continuous and scalar treatment variable on that unit’s outcome. Let $D_{1}$ (resp. $D_{2}$) denote the unit’s treatment at period 1 (resp. 2), and let $\mathcal{D}_{1}$ (resp. $\mathcal{D}_{2}$) be its support. Let $S=1\{D_{2}\neq D_{1}\}$ be an indicator equal to 1 if the unit’s treatment changes from period one to two, i.e. if they are a switcher.
For any $d\in\mathcal{D}_{1}\cup\mathcal{D}_{2}$, let $Y_{1}(d)$ and $Y_{2}(d)$ respectively denote the unit’s potential outcomes at periods 1 and 2 with treatment $d$, and let $Y_{1}$ and $Y_{2}$ denote their observed outcomes. Our potential outcome notation assumes that $Y_{1}$ does not depend on units’ period-two treatment, thus ruling out anticipation effects, a commonly-made assumption in the DID literature. Our notation also assumes that $Y_{2}$ does not depend on units’ period-one treatment, thus ruling out dynamic effects. This is not of essence in the two-periods case we focus on for now. This is of essence when we extend our estimators to the case where $T>2$, as we will explain in Section 5. Therein, we propose a modified version of our estimators, which partly alleviates this concern by allowing for dynamic effects up to a pre-specified treatment lag.
@@ 2.2 Assumptions
We make the following assumptions.
@@
(Parallel trends) $\forall$ $d_{1}\in\mathcal{D}_{1}$, $E(\Delta Y(d_{1})|D_{1}=d_{1},D_{2})=E(\Delta Y(d_{1})|D_{1}=d_{1})$.
Assumption 1 is a parallel-trends assumption, requiring that $\Delta Y(d_{1})$ be mean independent of $D_{2}$, conditional on $D_{1}=d_{1}$.
@@
(Bounded treatment and regularity of potential outcomes)
@@
    $\sup_{d_{1}\in\mathcal{D}_{1}}E[(\Delta Y(d_{1}))^{2}|D_{1}=d_{1}]<\infty$.
  3. 3.
    $\forall$ $t\in\{1,2\}$ and $\forall$ $(d,d^{\prime})\in\mathcal{D}_{t}^{2}$, there is a random variable $\overline{Y}\geq 0$ such that $|Y_{t}(d)-Y_{t}(d^{\prime})|\leq\overline{Y}|d-d^{\prime}|$, with $\sup_{(d_{1},d_{2})\in\text{Supp}(D_{1},D_{2})}E[\overline{Y}^{2}|D_{1}=d_{1},D_{2}=d_{2}]<\infty$.
Assumption 2 ensures that all the expectations below are well defined. It requires that the set of values that the period-one and period-two treatments can take be bounded. It also requires that the potential outcome functions be Lipschitz (with a unit-specific Lipschitz constant). This holds if $d\mapsto Y_{2}(d)$ is differentiable with respect to $d$ and has a bounded derivative.
@@
(iid sample) We observe $(Y_{i,1},Y_{i,2},D_{i,1},D_{i,2})_{1\leq i\leq n}$, which are independent and identically distributed vectors with the same probability distribution as $(Y_{1},Y_{2},D_{1},D_{2})$.